Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
capocasa, polini, sorrentino - 10:31 Saturday 12 June 2021 (52128) Print this report
Autocollimation of BAB in reflection from FCIM

In order to align FC with IR we wanted to auto-collimate its reflection from FCIM.

As a first thing we align FC for green:

1. We centered the camera in transmission with SQB2 M1 

2. We slightly adjusted FCIM to superpose the back reflection on the incoming beam

3. We had to move FCEM oplev setpoint quite a lot to find the flashes (TX -665 urad -> -716 urad, TY 2265 urad -> 2326 urad). To recover the alignment it was useful to check the camera looking at the cavity-side of FCIM.

Then we moved SQB1_M21 and SQB1_M22 to optimise and the beam that get back from the FI on SQB1, by looking it with a IR card. 

In order to better check the reflection and improve the optimisation, we spent a long time to align the mirrors at the output of the FI on SQB1 in order to send the reflection back on a camera and PD on EQB1. The path (which was very misalignmed) includes 2 steering mirror without picomotors (SQB1_M12 and SQB1_M14) and 2 small diaphragms. It difficult and extremely time-consuming to do it with the suspended bench as it requires to open the controls, move “blindly" the steering (the beam is not exiting the faraday with open control), close the controls, check the status and repeat again and again).

At the end, the reflection was reaching EQB1 but it is not well centred on the small diaphragm/optics. We should consider the possibility to block the bench for a better alignment and complete the autocollimation optimization.

Since the beam is coming out from the FI at a lower height, the first steering mirror (SQB1_M12) has to be much tilted to send it on the center of the second steering (SQB_M13). It seems that with the max vertical steering the beam is still very low on SQB1_M13.

Before leaving, we closed the doors of the SQB1 and SQB2 (left open in the same position for few days) and opened the controls, since changing the door position usually requires to recenter SBE.


In the morning, Fiodor restored the SubCarrier and removed some attenuators to increase the power. It reached 25 mW. The plan was to superpose it to the BAB and use it for the alignment of the FC but the automatic alignment system between SC and BAB was not working well and the two beams are visibly not well superposed already on SQB1. So we decided to proceed only with the BAB for the auto-collimation activity.

Images attached to this report
Comments to this report:
sorrentino - 12:40 Wednesday 16 June 2021 (52166) Print this report

With the recent tuning of the working point for the 3 QNR lasers, the temperature of SC laser crystal was decreased by only about 1 deg while the main and CC lasers were moved by about 1.5 deg. Thus the sign of the offset frequency of the SC laser from main laser was changed, and the PLL was locked at 1060 MHz instead of the nominal 1260 MHz. In order to engage the SC AA loop, which requires a frequency offset of 1260 MHz, on Friday 11/06 I changed the SC laser temperature from 30.26 deg to 29.41 deg, and I managed to engage the PLL at 1260 frequency offset. I could not reach the needed temperature with the control voltage from the DAC, and I observed that the control voltage saturates below DAC count = 22000 and above DAC count = 40000; the linear range of the control voltage is only from -1.6 V to +1.1 V. Thus I set the control voltage to DAC count = 32000 and I moved the laser temperature with the manual knob on the driver.

I engaged the SC AA loop, however the loop did not maximize the SC-CC beat. On 15/06 I verified that this was due to some clipping of the SC beam on the Faraday isolator, then I manually adjusted the aligment, I engaged the AA loop, and I verified that the beat note amplidute was maximized.

Search Help